# On Supportedness in Multi-Objective Combinatorial Optimization

### v1

@@ 1 Introduction
The set of efficient solutions and the set of non-dominated points are often decomposed into the sets of supported and unsupported efficient solutions and non-dominated points, respectively. Supported solutions are typically defined as solutions that can be obtained as optimal solutions of a weighted sum scalarization. The computation of unsupported non-dominated points requires different scalarization techniques and is often computationally more expensive. Moreover, for multi-objective combinatorial optimization problems (MOCO) the unsupported non-dominated points typically outnumber the supported ones, as observed in Visẽe et al., (1998). The determination of supported non-dominated points has gained attention for several reasons. First, these points are generally easier to determine than the unsupported non-dominated points. Second, they can serve as a foundation for the second phase of two-phase methods, which aims to generate the entire non-dominated point set using information derived from the supported non-dominated points (Pasternak and Passy, 1972; Visẽe et al., 1998; Hamacher et al., 2007; Przybylski et al., 2008; Przybylski et al., 2010b ; Eusébio and Figueira, 2009a ; Dai and Charkhgard, 2018,…). Recently, a computational study Sayin, (2024) showed that the set of supported non-dominated points can be used as a high-quality representation in multi-objective discrete optimization problems, focusing specifically on binary knapsack and assignment problems.
Several studies focus on identifying or analyzing the supported or extreme supported point set across various combinatorial problems, including bi- and multi-objective integer network flow problems (Eusébio and Figueira, 2009c, ; Medrano and Church,, 2015; Raith and Sedeño-Noda,, 2017; Könen and Stiglmayr, 2023a, ; Könen and Stiglmayr, 2023b, ), bi- and multi-objective minimum spanning trees (Sourd et al.,, 2006; da Silva and Clí­maco,, 2007; Correia et al.,, 2021), bi- and multi-objective shortest path problems (Edwin Romeijn and Smith,, 1999; Sedeño-Noda and Raith,, 2015; Medrano and Church,, 2014), bi- and multi-objective combinatorial unconstrained problems (Bökler,, 2018; Schulze et al.,, 2019), bi- and multi-objective knapsack problems (Visẽe et al.,, 1998; Argyris et al.,, 2011; Schulze,, 2017), and bi- and multi-objective assignment problems (Tuyttens et al.,, 2000; Gandibleux et al.,, 2003; Przybylski et al., 2010b, ). Additional work on general MOCO problems can be found in Gandibleux et al., (2001); Jesus, (2015); Sayin, (2024), among others.
However, despite their importance, several characterizations exist for supported efficient solutions and analogously for supported non-dominated points. These different definitions differ in the literature and sometimes even within a single publication. While these definitions are equivalent in the case of multi-objective linear problems (Isermann,, 1974; Ehrgott,, 2005), they can lead to different sets of supported efficient solutions and thus supported non-dominated point sets.
Supported non-dominated points for MOCO problems are often characterized as non-dominated points that lie on the boundary of the upper image and that they only lie on the non-dominated frontier, whereas unsupported solutions are characterized as non-dominated points that lie in the interior of the upper image (Eusébio and Figueira, 2009c, ; Przybylski et al., 2010a, ).
However, depending on the definition, unsupported non-dominated points may exist that lie on the boundary of the upper image or supported non-dominated points that lie on weakly non-dominated faces and thus do not lie on the non-dominated frontier. In this case, they cannot be obtained as optimal solutions of a weighted sum problem with weights strictly greater than zero.
This motivates us to distinguish between supported efficient solutions and weakly supported efficient solutions. An efficient solution is denoted as weakly supported if it is an optimal solution of a weighted sum scalarization with non-negative weights. In contrast to the definition of supportedness, weakly supportedness allows single weights to have a value of zero. In order to avoid confusion, note that weakly supported efficient solutions are efficient. However, their images may lie on weakly non-dominated faces. An unsupported efficient solution is then defined as an efficient solution that is neither supported nor weakly supported. Then the following characterizations for MOCO problems hold: While weakly supported non-dominated points lie on the boundary of the upper image, supported non-dominated points lie only on the non-dominated frontier, i. e., only on maximally non-dominated faces. The unsupported non-dominated points lie in the interior of the upper image. We will present an example where the set of supported efficient solutions is a proper subset of all weakly efficient solutions in Section 3. The clear distinction between the sets of supported non-dominated and weakly supported non-dominated solutions is also necessary as the corresponding problems may differ in their output time complexity. In particular, in the case of the minimum cost flow problem, it can be shown that supported efficient solutions can be determined in output-polynomial time, whereas this is not the case for the weakly supported solutions unless $\mathbf{P}=\mathbf{NP}$ (Könen and Stiglmayr, 2023a, ). Note that the computation of the weakly supported solutions is as “hard” as the determination of all non-dominated points.
While the above characterization of supported points applies to MOCO problems, they do not extend to general MOO problems, as illustrated in Example 3.14 in Ehrgott, (2005). This distinction arises because the weighted sum scalarization method with weights strictly positive is only capable of identifying properly supported non-dominated points in the sense of Geoffrion (Geoffrion,, 1968), as discussed in Ehrgott, (2005). For MOCO problems, the distinction between efficient and properly efficient solutions vanishes, since every efficient solution is also properly efficient (Ehrgott,, 2005). However, for general MOO problems, the set of properly supported non-dominated points may form a strict subset of the non-dominated points on the non-dominated frontier, highlighting the limitations of the weighted sum approach in capturing the entire set of supported solutions in general MOO problems (Ehrgott,, 2005).
This work focuses on the supportedness in multi-objective discrete problems. However, further investigation into the supportedness of general MOO problems remains sparse and presents an important area for future research. A recent presentation at the EURO 2024 conference (Chlumsky-Harttmann and Schöbel,, 2024) provided a first approach towards a categorization of supportedness definitions for general MOO problems.
The remainder of the paper is structured as follows. Section 2 presents notations and preliminaries. In Section 3 the different definitions and characterizations found in the literature are presented. A counterexample is given that shows that these definitions are not equivalent for MOCO or general non-convex multi-objective optimization problems. Additionally, it discusses the distinction between supported and weakly supported non-dominated points. Finally, Section 4 concludes the findings and outlines directions for future research.
2 Multi-Objective Optimization and Combinatorial Optimization
This section formally introduces the multi-objective combinatorial optimization problem along with some of the most important results and properties of general multi-objective optimization. For a comprehensive introduction to multi-objective optimization, we refer to Steuer, (1986) and Ehrgott, (2005).
A multi-objective optimization (MOO) problem is defined as
@@ 2 Multi-Objective Optimization and Combinatorial Optimization
where $f\colon\mathds{R}^{n}\rightarrow\mathds{R}^{p}$ is a vector-valued objective function composed of $p\geq 2$ real-valued objective functions $f_{k}\colon\mathds{R}^{n}\rightarrow\mathds{R}$ for $k\in\{1,\ldots,p\}$ and $\mathcal{X}\subseteq\mathds{R}^{n}$ denotes the set of feasible solutions. We call $\mathds{R}^{n}$ the decision space and $\mathds{R}^{p}$ the objective space of (MOO). The image of the feasible set $\mathcal{Y}\coloneqq f(\mathcal{X})=\{f(x)\colon x\in\mathcal{X}\}\subseteq\mathds{R}^{p}$ is called the set of feasible outcome vectors in the objective space.
We assume that the objective functions are conflicting, which implies that we exclude the existence of an ideal solution that minimizes all objectives simultaneously. We use the Pareto concept of optimality, which is based on the component-wise order in $\mathds{R}^{p}$. Let $y^{1},y^{2}\in\mathbb{R}^{p}$. We write:
$y^{1}\leqslant y^{2}\iff y_{k}^{1}\leqq y^{2}_{k}\quad\forall k\in\{1,\dots,p\}\quad\text{ and }\quad y^{1}\neq y^{2}$
The non-negative orthant of $\mathds{R}^{p}$ is denoted by $\mathbb{R}_{\geqq}^{p}:=\left\{x\in\mathbb{R}^{p}:x\geqq 0\right\}$ and analogously $\mathds{R}_{\geq}^{p}$ and its interior $\mathbb{R}_{>}^{p}$. In our notation, we also use the Minkowski sum and the Minkowski product of two sets $A,B\subseteq\mathds{R}^{d}$, which are defined as $A+B\coloneqq\{a+b\colon a\in A,b\in B\}$ and $A\cdot B\coloneqq\{a\cdot b\colon$ $a\in A,b\in B\}$, respectively. For a closed set $C$, we define the boundary of $C$ by $\partial C$. The polyhedron $\mathcal{Y}^{\geqq}:=\operatorname{conv}(\mathcal{Y}_{N}+\mathds{R}^{p}_{\geqq})$ is called the upper image of $\mathcal{Y}$.
A specific, well-studied class of MOO are multi-objective linear programs (MOLP), which are given by
with $\mathcal{X}\coloneqq\{x\in\mathds{R}^{n}\colon A\,x\leq b\}$. Thereby, the rows $c^{k}$ ($k=1,\ldots,p$) of the cost matrix $C\in\mathbb{R}^{p\times n}$ contain the coefficients of the $p$ linear objective functions $y_{k}(x)=c^{k}\,x$ for $k\in\{1,\ldots,p\}$ and the matrix $A\in\mathbb{R}^{m\times n},b\in\mathbb{R}^{m}$ formulates the $m$ linear constraints.
@@
A feasible solution $x^{*}\in\mathcal{X}$ is called efficient if there is no feasible solution $x\in\mathcal{X}$ such that $f(x)\leq f\left(x^{*}\right)$. If $x^{*}$ is efficient, then its image $f\left(x^{*}\right)$ is called non-dominated point. The set of efficient solutions is denoted by $\mathcal{X}_{E}\subseteq\mathcal{X}$ and the set of non-dominated points by $\mathcal{Y}_{N}\subseteq\mathcal{Y}$. Moreover, a feasible solution $\hat{x}$ is called weakly efficient if there is no other solution $x\in\mathcal{X}$ such that $f(x)<f(\hat{x})$.
A feasible solution $x^{*}\in\mathcal{X}$ is called properly efficient in Geoffrion’s sense (Geoffrion,, 1968), if it is efficient and if there is a real number $M>0$ such that for all $k\in\{1,\ldots,p\}$ and $x\in\mathcal{X}$ satisfying $f_{k}(x)<f_{k}({x}^{*})$ there exists an index $j$ such that $f_{j}({x}^{*})<f_{j}(x)$ such that
@@
In the following, we assume that the multi-objective optimization problem has at least one non-dominated point, i.e., $\mathcal{Y}_{N}\neq\varnothing$. Scalarization approaches are the most common solution techniques in multi-objective optimization, which rely on replacing the multi-objective problem by a sequence of single-criteria optimization problems. In the weighted-sum scalarization a convex combination (with weights $\lambda_{i}$) of the objective functions is optimized over the feasible set $\mathcal{X}$. Then, the set of normalized weighting vectors (or weight space) is defined as the set $\Lambda_{p}=\{\lambda\in\mathds{R}^{p}_{>}\colon\sum_{i=1}^{p}\lambda_{i}=1\}$ or $\Lambda_{p}^{0}=\{\lambda\in\mathds{R}^{p}_{\geq}\colon\sum_{i=1}^{p}\lambda_{i}=1\}$ if weights equal to zero are included.
@@
If $\lambda\in\Lambda_{p}^{0}$, every optimal solution of $P_{\lambda}$ is a weakly efficient solution of (MOO). Moreover, every optimal solution of $P_{\lambda}$ problem is (properly) efficient for (MOO), if $\lambda\in\Lambda_{p}$.
@@
Hence, for a MOLP, the efficient set $\mathcal{X}_{E}$ and the union over all sets of optimal solutions of $P_{\lambda}$ for $\lambda\in\Lambda_{p}$ are identical.
However, in the non-linear case and for multi-objective combinatorial (MOCO) problems, non-dominated points can be located also in the interior of the upper image, i.e., $\operatorname{int}(\operatorname{conv}(\mathcal{Y}+\mathds{R}^{p}_{\geqq}))$ and cannot be determined using a weighted-sum method.
@@
where $\mathcal{X}\coloneqq\{x\in\{0,1\}^{n}\colon Ax=b\}$, with the cost matrix $C\in\mathbb{Z}^{p\times n}$ containing the rows $c^{k}$ of coefficients of $p$ linear objective functions $y_{k}(x)=c^{k}\,x$ for $k\in\{1,\ldots,p\}$ and $A\in\mathbb{Z}^{m\times n},b\in\mathbb{Z}^{m}$ describing the $m$ constraints. The constraints define combinatorial structures such as paths, trees, or cycles in a network or partitions of a set. All coefficients are assumed to be integers.
This definition includes various problems such as multi-objective spanning trees, shortest paths, network flow, knapsack, and assignment problems.
For MOLP or MOCO problems, the non-dominated frontier can be defined in the following way:
@@
The non-dominated frontier is the set
The non-dominated frontier equals the non-dominated set of $\operatorname{conv}(\mathcal{Y}_{N})$ and can be characterized as the set of the maximal non-dominated faces of the upper image $\operatorname{conv}(\mathcal{Y})$ (Ehrgott,, 2005). Hereby, a non-dominated face $F\subseteq\mathcal{Y}^{\geqq}$ is a face of the upper image $\mathcal{Y}^{\geqq}$ such that all its points are non-dominated with respect to $\mathcal{Y}^{\geqq}$. A face $F$ is called maximally non-dominated if there is no other non-dominated face $G$ of $\mathcal{Y}^{\geqq}$ such that $F$ is a proper subset of $G$. This would imply that the dimension of $G$ is be greater than the dimension of $F$.
The preimage $F_{\mathcal{X}}$ of a maximally non-dominated face $F_{\mathcal{Y}}$ of $\mathcal{Y}^{\geqq}$, i.e., all solutions whose image lies in $F_{\mathcal{Y}}$, is denoted as the maximally efficient face. Note that $F_{\mathcal{X}}$ might not be a face of the feasible set, the polyhedron $X$, since multiple feasible solutions (located on different faces of $\mathcal{X}$) could map to the same non-dominated point $y\in\mathcal{Y}$.
Two major challenges are associated with MOCO problems. Firstly, they are theoretically more challenging than their single-objective counterparts, falling into the class of computationally intractable problems (Ehrgott,, 2005). Secondly, in MOCO or MOILP problems, non-dominated points located in the interior of the upper image can occur, which typically outnumber the set of solutions on the non-dominated frontier (see, e.g., Visẽe et al., 1998). Hence, a distinction between different classes of efficient solutions is required. An intuitive definition of supportedness is the following.
@@
    An efficient solution is called supported efficient solution if it is an optimal solution to the weighted sum scalarization $P_{\lambda}$ for $\lambda\in\Lambda_{d}$, i.e., an optimal solution to a single objective weighted-sum problem where the weights are strictly positive. Its image is called supported non-dominated point; we use the notation $\mathcal{X}_{S}$ and $\mathcal{Y}_{S}$. For discrete problems, supported non-dominated points are located on the non-dominated frontier, i.e., located on the union of the maximal non-dominated faces of the upper image.
@@
    Extreme supported solutions are those solutions whose image lies on the vertex set of the upper image. Their image is called an extreme supported non-dominated point.
@@
Figure 1 illustrates supported extreme, supported, and unsupported non-dominated points, as well as the upper image in the bi-objective case.
$y_{1}$$y_{2}$$\operatorname{conv}(\mathcal{Y}+\mathds{R}^{2}_{\geqq})$extreme supported non-dominatedsupported non-dominatedunsupported non-dominateddominated
@@
A recent study by Sayin, (2024) shows that the set of (extreme) supported non-dominated points is a good representations of the non-dominated set. Therefore, one is interested in solution methods that can determine supported or extreme supported efficient solutions.
However, in the context of MOCO problems, other definitions and characterizations for supportedness are also used. The following chapter proves that these definitions are not equivalent in the case of discrete or more general optimization problems. As a result, they generate different sets of supported efficient solutions and, consequently, different supported non-dominated point sets with different properties.

### v2

@@ 1 Introduction
Multi-objective optimization (MOO) is concerned with optimizing multiple conflicting objectives simultaneously. For a comprehensive introduction to multi-objective optimization, we refer to Steuer, (1986) and Ehrgott, (2005).
Definition 1.1
A multi-objective optimization problem (MOO) is defined as
@@
where $f\colon\mathds{R}^{n}\rightarrow\mathds{R}^{p}$ is a vector-valued objective function composed of $p\geq 2$ real-valued objective functions $f_{k}\colon\mathds{R}^{n}\rightarrow\mathds{R}$ for $k\in\{1,\ldots,p\}$ and $\mathcal{X}\subseteq\mathds{R}^{n}$ denotes the set of feasible solutions. We call $\mathds{R}^{n}$ the decision space and $\mathds{R}^{p}$ the objective space. The image of the feasible set
$\mathcal{Y}\coloneqq f(\mathcal{X})=\{f(x)\colon x\in\mathcal{X}\}\subseteq\mathds{R}^{p}$
is called the set of feasible outcome points in objective space.
The non-negative orthant of $\mathds{R}^{p}$ is denoted by $\mathbb{R}_{\geqq}^{p}:=\left\{x\in\mathbb{R}^{p}:x\geqq 0\right\}$ and analogously its interior $\mathds{R}_{>}^{p}$ and $\mathds{R}_{\geq}^{p}:=\left\{x\in\mathds{R}^{p}:x\geqq 0,x\neq 0\right\}$. In our notation, we also use the Minkowski sum and the Minkowski product of two sets $A,B\subseteq\mathds{R}^{p}$, which are defined as $A+B\coloneqq\{a+b\colon a\in A,b\in B\}$ and $A\cdot B\coloneqq\{a\cdot b\colon$ $a\in A,b\in B\}$, respectively. For a closed set $C$, we define the boundary of $C$ by $\partial C$.
The objective functions are typically assumed to be conflicting, implying the absence of an ideal solution that minimizes all objectives simultaneously. Instead, we rely on the Pareto concept of optimality, based on the component-wise order in $\mathds{R}^{p}$.
@@
Let $y^{1},y^{2}\in\mathbb{R}^{p}$. We write:
  • $y^{1}\leqq y^{2}$ if $y_{i}^{1}\leq y^{2}_{i}$ for $i=1,\ldots,p$,
  • $y^{1}\leq y^{2}$ if $y^{1}\leqq y^{2}$ but $y^{1}\neq y^{2}$ and
  • $y^{1}<y^{2}$ if $y_{i}^{1}<y^{2}_{i}$ for $i=1,\ldots,p$.
Definition 1.3
A feasible solution $x^{*}\in\mathcal{X}$ is called efficient if there is no $x\in\mathcal{X}$ such that $f(x)\leq f\left(x^{*}\right)$. The image $f(x^{*})$ is then called a non-dominated point. The set of efficient solutions is denoted by $\mathcal{X}_{E}\subseteq\mathcal{X}$ and the set of non-dominated points by $\mathcal{Y}_{N}\subseteq\mathcal{Y}$.
A feasible solution $\hat{x}\in X$ is called weakly efficient if no $x\in\mathcal{X}$ exists with f(x)<f(x^)f(x)<f($\hat${x}). The image f(x^)f($\hat${x}) is then called a weakly non-dominated point. Furthermore, a feasible solution $x^{*}\in\mathcal{X}$ is called properly efficient (in Geoffrion’s sense (Geoffrion,, 1968)), if it is efficient and if there is a real number $M>0$ such that for all $k\in\{1,\ldots,p\}$ and $x\in\mathcal{X}$ satisfying $f_{k}(x)<f_{k}({x}^{*})$ there exists an index $j$ such that $f_{j}({x}^{*})<f_{j}(x)$ such that
@@
In the following, we assume that the multi-objective optimization problem has at least one non-dominated point, i.e., $\mathcal{Y}_{N}\neq\varnothing$. The polyhedron $\mathcal{Y}^{\geqq}:=\operatorname{conv}(\mathcal{Y}_{N}+\mathds{R}^{p}_{\geqq})$ is called the upper image of $\mathcal{Y}$.
Scalarization approaches are the most common solution techniques in multi-objective optimization, which rely on replacing the multi-objective problem by a sequence of single-criteria optimization problems. In the weighted-sum scalarization a convex combination (with weights $\lambda_{i}$) of the objective functions is optimized over the feasible set $\mathcal{X}$. Then, the set of normalized weighting vectors (or weight space) is defined as the set $\Lambda_{p}=\{\lambda\in\mathds{R}^{p}_{>}\colon\sum_{i=1}^{p}\lambda_{i}=1\}$ or $\Lambda_{p}^{0}=\{\lambda\in\mathds{R}^{p}_{\geq}\colon\sum_{i=1}^{p}\lambda_{i}=1\}$ if weights equal to zero are included.
@@
If $\lambda\in\Lambda_{p}^{0}$, every optimal solution of $P_{\lambda}$ is a weakly efficient solution of (MOO). Moreover, every optimal solution of $P_{\lambda}$ problem is properly efficient for (MOO), if $\lambda\in\Lambda_{p}$.
A specific, well-studied class of MOO are multi-objective linear programs (MOLP) and multi-objective integer linear programs (MOILP).
Definition 1.6
A multi-objective linear optimization problem (MOLP) is defined by
with $\mathcal{X}\coloneqq\{x\in\mathds{R}^{n}\colon Ax\leqq b,x\geqq 0\}$. Thereby, the rows $c^{k}$ ($k=1,\ldots,p$) of the cost matrix $C\in\mathbb{R}^{p\times n}$ contain the coefficients of the $p$ linear objective functions $y_{k}(x)=c^{k}\,x$ for $k\in\{1,\ldots,p\}$ and the matrix $A\in\mathbb{R}^{m\times n},b\in\mathbb{R}^{m}$ formulates the $m$ linear constraints.
Definition 1.7
If all decision variables are further restricted to the set of integers, i.e., $x\in\mathds{Z}$, we obtain a multi-objective integer linear problem (MOILP).
$\min_{x\in\mathcal{X}}y(x)=\left(y_{1}(x),\ldots,y_{p}(x)\right)^{\top}=Cx,$ (MOILP)
with $\mathcal{X}\coloneqq\{x\in\mathds{Z}^{n}\colon Ax\leqq b,x\geqq 0\}$ and all other elements as described in (MOLP).
@@
Hence, for a MOLP, the efficient set $\mathcal{X}_{E}$ and the union over all sets of optimal solutions of $P_{\lambda}$ for $\lambda\in\Lambda_{p}$ are identical. Thus, in multi-objective linear programs (MOLP) every non-dominated point lies on the non-dominated frontier, which can be defined in the following way.
Definition 1.9 (Hamacher et al., (2007))
The non-dominated frontier is the set
Note that Hamacher uses the wording efficient frontier instead of the non-dominated frontier. The non-dominated frontier equals the non-dominated set of $\operatorname{conv}(\mathcal{Y}_{N})$ and can be characterized as the set of the maximal non-dominated faces of the upper image $\mathcal{Y}^{\geqq}$ (Ehrgott,, 2005). Hereby, a non-dominated face $F\subseteq\mathcal{Y}^{\geqq}$ is a face of the upper image $\mathcal{Y}^{\geqq}$ such that every point on $F$ is non-dominated with respect to $\mathcal{Y}^{\geqq}$. Similarly, $F\subseteq\mathcal{Y}^{\geqq}$ is called a weakly non-dominated face if all its points are weakly non-dominated. A face $F$ is called maximally non-dominated if there is no other non-dominated face $G$ of $\mathcal{Y}^{\geqq}$ such that $F$ is a proper subset of $G$. This would imply that the dimension of $G$ is greater than the dimension of $F$. The preimage $F_{\mathcal{X}}$ of a maximally non-dominated face $F_{\mathcal{Y}}$ of $\mathcal{Y}^{\geqq}$, i.e., all solutions whose image lies in $F_{\mathcal{Y}}$, is denoted as the maximally efficient face. Note that $F_{\mathcal{X}}$ might not be a face of the feasible set, the polyhedron $\mathcal{X}$, since multiple feasible solutions (located on different faces of $\mathcal{X}$) could map to the same non-dominated point $y\in\mathcal{Y}$.
$y_{3}$$y_{1}$$y_{2}$
Figure 1: Illustration of dominance relations in faces: the red triangular face represents a maximal non-dominated face; its red boundary edges are non-dominated faces; and the light blue faces are weakly non-dominated.
Figure 1 illustrates the geometric structure of non-dominated and weakly non-dominated faces for a continuous multi-objective minimization problem, where the feasible region consists of a cube with a cut at its lower-left front corner. The red triangular region corresponds to a maximal non-dominated face, meaning no point in this region is dominated by another feasible point, and this face cannot be extended to a larger non-dominated face within the feasible set. The bold red edges of this triangle represent lower-dimensional non-dominated faces, while the light-blue shaded cube faces indicate weakly non-dominated faces, which are optimal in some but not all directions.
In contrast to multi-objective linear programs (MOLP), where every non-dominated point lies on the non-dominated frontier, and thus on the convex hull of the upper image, MOILPs or MOCO problems may contain non-dominated points that are located in the interior of the upper image, i.e., in $\operatorname{int}(\operatorname{conv}(\mathcal{Y}_{N}+\mathds{R}^{p}_{\geqq}))$ and cannot be generated through any weighted-sum scalarization.
@@
where $\mathcal{X}\coloneqq\{x\in\{0,1\}^{n}\colon Ax=b\}$, with the cost matrix $C\in\mathbb{Z}^{p\times n}$ containing the rows $c^{k}$ of coefficients of $p$ linear objective functions $y_{k}(x)=c^{k}\,x$ for $k\in\{1,\ldots,p\}$ and $A\in\mathbb{Z}^{m\times n},b\in\mathbb{Z}^{m}$ describing the $m$ constraints. The constraints define combinatorial structures such as paths, trees, or cycles in a network or partitions of a set.
This definition includes various problems such as multi-objective spanning trees, shortest paths, knapsack and assignment problems.
Two major challenges are associated with MOCO problems. Firstly, they are theoretically more challenging than their single-objective counterparts, falling into the class of computationally intractable problems (Ehrgott,, 2005). Secondly, in MOCO problems, non-dominated points located in the interior of the upper image can occur, which typically outnumber the set of solutions on the non-dominated frontier (see, e.g., Visẽe et al., 1998). Hence, a distinction between different classes of efficient solutions is required. An intuitive definition of supportedness is the following.
@@
    An efficient solution is called supported efficient solution if it is an optimal solution to the weighted sum scalarization $P_{\lambda}$ for $\lambda\in\Lambda_{p}$, i.e., an optimal solution to a single objective weighted-sum problem where the weights are strictly positive. Its image is called supported non-dominated point; we use the notation $\mathcal{X}_{S}$ and $\mathcal{Y}_{S}$ for the supported efficient solution set and supported non-dominated point set, respectively. For discrete problems, supported non-dominated points are located on the non-dominated frontier, i.e., located on the union of the maximal non-dominated faces of the upper image.
@@
    Extreme supported efficient solutions are those solutions whose image lies on the vertex set of the upper image. Their image is called an extreme supported non-dominated point.
@@
Figure˜2 illustrates supported extreme, supported, and unsupported non-dominated points, as well as the upper image in the bi-objective case.
$y_{1}$$y_{2}$$\operatorname{conv}(\mathcal{Y}_{N}+\mathds{R}^{2}_{\geqq})$extreme supported non-dominatedsupported non-dominatedunsupported non-dominateddominated
@@
As unsupported non-dominated points cannot be computed through a weighted sum scalarization, different scalarization techniques are required that are often computationally more expensive.
The determination of supported non-dominated points has gained attention for several reasons. First, these points are generally easier to determine than the unsupported non-dominated points and, moreover, the unsupported non-dominated points in MOCO typically outnumber the supported ones, as observed in Studies as Visẽe et al., (1998).
Second, they can serve as a foundation for the second phase of two-phase methods, which aims to generate the entire non-dominated point set using information derived from the supported non-dominated points (Pasternak and Passy, 1972; Visẽe et al., 1998; Hamacher et al., 2007; Przybylski et al., 2008; Przybylski et al., 2010b ; Eusébio and Figueira, 2009a ; Dai and Charkhgard, 2018,…). Recently, a computational study Sayin, (2024) showed that the set of supported non-dominated points can be used as a high-quality representation in multi-objective discrete optimization problems, focusing specifically on binary knapsack and assignment problems.
Several studies focus on identifying or analyzing the supported or extreme supported point set across various combinatorial problems, including bi- and multi-objective integer network flow problems (Eusébio and Figueira, 2009c, ; Medrano and Church,, 2015; Raith and Sedeño-Noda,, 2017; Könen and Stiglmayr,, 2025, 2023), bi- and multi-objective minimum spanning trees (Sourd et al.,, 2006; da Silva and Clí­maco,, 2007; Correia et al.,, 2021), bi- and multi-objective shortest path problems (Edwin Romeijn and Smith,, 1999; Sedeño-Noda and Raith,, 2015; Medrano and Church,, 2014), bi- and multi-objective combinatorial unconstrained problems (Bökler,, 2018; Schulze et al.,, 2019), bi- and multi-objective knapsack problems (Visẽe et al.,, 1998; Argyris et al.,, 2011; Schulze,, 2017), and bi- and multi-objective assignment problems (Tuyttens et al.,, 2000; Gandibleux et al.,, 2003; Przybylski et al., 2010b, ). Additional work on general MOCO problems can be found in Gandibleux et al., (2001); Jesus, (2015); Sayin, (2024), among others.
However, despite their significance, several alternative definitions (as in Definition˜1.11) and characterizations for supportedness are used. These different definitions differ in the literature and sometimes even within a single publication. The following chapter proves that these definitions are not equivalent in the case of discrete or more general optimization problems. As a result, they generate different sets of supported efficient solutions and, consequently, different supported non-dominated point sets with varying properties.
This motivates us to distinguish between supported efficient solutions and weakly supported efficient solutions. An efficient solution is denoted as weakly supported if it is an optimal solution of a weighted sum scalarization with non-negative weights. In contrast to the definition of supportedness, weakly supportedness allows single weights to have a value of zero. Note that the terminology weakly supported efficient solutions, despite the fact that these solutions are still efficient, is motivated by geometric considerations: any weakly supported non-dominated point that is not a supported non-dominated point lies on a weakly non-dominated face, but not on a non-dominated face. In contrast, supported points necessarily lie on non-dominated faces.
This work focuses on the theoretical understanding of supportedness in MOCO problems. In particular, we analyze the structural properties of supported and weakly supported non-dominated points and highlight important distinctions through counterexamples. However, further investigation into the supportedness of general MOO problems remains sparse and presents an important area for future research. A recent study in (Chlumsky-Harttmann,, 2025) provided a first approach toward a categorization of supportedness definitions for general MOO problems.
The remainder of the paper is structured as follows. In Section˜2 the different definitions and characterizations of supportedness found in the literature are presented. A counterexample is given that shows that these definitions are not equivalent for MOILPs and MOCO problems. Additionally, it discusses the distinction between supported and weakly supported non-dominated points. Finally, Section˜3 concludes the findings and outlines directions for future research.